Use Excel to calculate for the following dataset:

X Y
1 10
1 12
3 13
3 17
5 17
5 21

the correlation coefficient:   Please round to two decimals

the value of the intercept, b0:  

the value of the slope, b1:    

the predicted value of Y when X =5:    
Answer #1

Correlation coefficient: 0.8835

Intercept,b0: 9

Slope, b1: 2

Least square linear regression is given by,

ŷ = b0+b1x

ŷ = 9+2x

We have to predict y for x = 5

ŷ = 9+2*5

ŷ = 19